Which does not show paramagnetismClo2No2Clo2^-No

Grade:12

1 Answers

Mahima Kanawat
6 years ago
Dear student ....To find paramgnetism count the valence electrons in elements of compound and add them , if they are odd then it is paramagnetic...ClO2 = 7+6+6=19; NO2 = 5+6+6=17 ; ClO2- =7+6+6+1= 20 ; NO =5+6 = 11..Therefore ,ClO2- is not paramagnetic ....With Regards ....Mahima....(Askiitians Forum Expert )